Ricardo is 4 years less than twice as old as his sister. The sum of their ages is 20. How old is Ricardo?
Answer by gsmani_iyer(178)



Let sister's age = x

So, Ricardo's age = 2x-4 (Given)

The sum of the ages = x + 2x - 4 = 20 (Given)

= 3x - 4 = 20

= 3x = 20 + 4 = 24

= x = 24/3 = 8.

So the sister's age = 8 yrs.
and Ricardo's age = (2*8)-4 = 16-4 = 12 yrs. Answer.